# Kevin Lin3, Xiaofei Wang3, Zhengyuan Yang3, Linjie Li3, Chung-Ching Lin3, Jianwen Xie4

### v1

@@ 5.4.1 Aesthetics
5.4.2 Low-level image statistics
@@
5.5 Multi-turn editing vs. complex editing
We compare two strategies for applying multiple edits. In multi-turn editing, instructions are executed sequentially: apply instruction 1, then apply instruction 2 to the result, and so on. In complex editing, we concatenate $C$ instructions into a single prompt and perform one edit (“complex level $C$, $C\in\{1,2,3\}$”). We evaluate two metrics: (i) the Turn-3 instruction-following rate (all edits up to turn 3 succeed), and (ii) the marginal success of the final instruction in a $C$-instruction complex prompt.
Empirically (Table 4; Figure 11), flow-matching editors (short-history condition) behave differently from AR editors when comparing multi-turn and complex modes. AR models tend to achieve higher final-task success in the multi-turn setting, demonstrating the benefit of a step-by-step “chain of edits” (analogous to chain-of-thought for reasoning). By contrast, many flow-matching models obtain better final-task performance under complex (single-shot) editing: sequential edits amplify failure modes (most notably exposure-related drift), which degrades multi-turn performance.
Practical recommendation: For true step-by-step workflows, we recommend using autoregressive architectures. For flow-matching editors, it is preferable to either (i) pack multiple edits into a single complex prompt when the instruction sequence is known in advance, or (ii) increase the model’s history window and/or add inter-turn regularization (e.g., histogram matching, exposure penalties, or periodic reconstruction passes) to mitigate cumulative drift.
Technique Model Multi-turn (T3) Complex (C3)
Autoregressive Nano Banana 35.35 28.14
GPT-Image-1 38.35 28.78
Gemini 2.0 Flash 28.42 21.89
Flow Matching Qwen-Image-Edit 22.55 27.62
Step1X-Edit 17.83 15.73
FLUX.1-Kontext-dev 16.61 19.58
OmniGen 10.66 11.01
Diffusion AnyEdit 7.22 2.80
UltraEdit 6.36 8.22
MagicBrush 4.90 4.55
IP2P 2.80 2.80
6 Limitation and Discussion
Given the object-centric evaluations conducted in this study, several limitations warrant consideration. First, our instruction types are limited to object-centric prompts, which may not capture the full range of creative editing requests typical in real-world scenarios. Future research should explore a broader spectrum of instructions, including those involving stylistic changes or complex narrative elements.
Additionally, while our work provides a reliable and comprehensive evaluation framework for multi-turn editing, it does not apply the evaluation results to improve the editing models themselves. A straightforward extension would be to use evaluation scores for Best-of-N selection to improve inference-time performance. Future work could also explore post-training methods such as reinforcement learning, treating the evaluation scores as reward signals.
7 Conclusion
We introduced EdiVal-Agent, a scalable, automated, and interpretable framework for evaluating instruction-based image editing. By leveraging symbolic object decomposition, structured instruction generation, and a hybrid evaluation pipeline integrating both specialist tools and vision-language reasoning models, EdiVal-Agent enables fine-grained, object-centric assessment of modern editing systems. Our design emphasizes transparency, extensibility, and real-world applicability—providing a new standard for evaluating multi-turn, compositional visual editing. We hope EdiVal-Agent will serve as a valuable resource for benchmarking, diagnosing, and advancing the next generation of instruction-based editing models.
Appendix A Algorithmic Details
@@ A.1 Decomposition
  • Object listing: We use GPT-4o with the prompt in Section A.2.1. The model returns a JSON with one entry per object and a terminal aggregated string key ‘‘All Objects’’.
@@ A.2 Instruction Generation
8:  if not feasible($\tau_{t},\mathcal{O}$then $\tau_{t}\leftarrow\texttt{subject\_add}$
@@
10:  $F_{t}\leftarrow\text{format\_instruction}(\tau_{t},\mathcal{O})$ $\triangleright$ Query VLM by prompts in Section A.2.2
@@
A.2.1 Object Listing Prompt
@@ A.3 Evaluation
We report HPSv3 Ma et al. (2025a) plausibility and aesthetics, plus luminance metrics. Quality is not folded into the Overall score.
@@
A.4 Overall Score and Aggregation Details
@@ A.5 Model generations
Prompts are provided in full (Section A.2); thresholds are specified above. Grounding uses SwinT-OGC weights; consistency uses DINOv3 ViT-B/16; the quality head follows our RAHF implementation, and HPSv3 is included when available. All other parameters are left at defaults.
Appendix B Additional Evaluation Results
@@
For the multi-turn editing mode, the overall instruction-following success rate is reported in Table 5, while success rates for individual instruction types appear in Tables 6 and 7. Consistency results are summarized in Table 11. We also observed that some input images are non-square after resizing, which can leave black padding on the top/bottom or left/right edges. Certain editing models, such as GPT-Image-1 and Qwen-Image-Edit, attempt to fill these areas, whereas others preserve them. To account for this, we separately report consistency for square (Table 12) and non-square inputs (Table 13). The conclusions remain consistent with the overall evaluation. Quality results for multi-turn editing are presented in Table 15.
For the complex editing mode, the overall instruction-following success rate is shown in Table 8, and per-instruction-type results are in Tables 9 and 10. Consistency and quality results are reported in Tables 14 and 16, respectively.
@@
GPT-4o 73.12 54.89 38.35 73.12 74.44 73.12
Nano 70.70 50.66 35.35 70.70 72.59 68.24
Gemini 68.07 45.96 28.42 68.07 67.72 68.42
QWEN 72.90 44.06 22.55 72.90 62.94 56.12
StepEdit 61.89 34.97 17.83 61.89 59.09 53.32
FLUX 59.97 32.69 16.61 59.97 56.29 51.40
@@
GPT4o 73.62 71.27 67.27 90.00 85.72 82.16 80.33 78.80 79.88 85.69 81.86 80.85
nano 88.68 85.19 83.31 96.88 93.79 92.41 89.73 87.64 88.48 94.21 93.77 93.36
Gemini 84.16 79.91 71.13 91.50 91.83 87.07 88.98 87.07 85.13 95.19 93.35 93.34
QWEN 75.36 72.68 68.85 88.50 88.98 82.71 81.66 78.47 77.04 91.16 87.85 85.78
StepEdit 87.40 84.03 78.27 97.43 92.42 88.15 93.14 91.57 89.40 98.10 97.12 96.01
FLUX 92.55 88.92 84.04 96.49 93.57 92.19 92.83 90.61 88.34 96.91 95.74 94.73
@@
GPT4o 73.22 68.53 67.20 88.60 86.88 83.99 79.57 78.27 77.35 77.20 75.91 75.14
nano 90.34 85.26 84.50 97.74 95.94 94.88 92.69 90.29 89.17 96.16 94.81 93.95
Gemini 85.69 76.79 72.12 96.14 93.23 90.10 90.91 86.23 83.99 95.02 93.12 91.51
QWEN 77.32 71.43 68.47 91.65 89.53 88.10 83.68 79.23 76.23 83.78 80.37 77.43
StepEdit 88.26 81.37 77.22 97.33 95.77 93.77 94.01 90.53 88.73 98.26 97.09 95.70
FLUX 92.67 87.80 85.44 98.15 96.92 95.54 94.57 91.70 90.10 96.30 94.98 94.05
@@
Appendix C Human Agreement
@@
We conducted human study on edits made by four exemplary models—Step1X-Edit, AnyEdit, Gemini-Flash 2.0, and Flux.1-Kontext-dev—on EdiVal-Bench, generated by EdiVal-Agent as described in Section A.2. For each edit, we collected two human ratings, yielding a total of $572\times 4\times 2=4{,}576$ annotations. Depending on the prompt (which affected the editing instruction), each annotation took about 1–2 minutes. Raters were recruited online, each holding at least a bachelor’s degree. They were shown the original image, the edited image, and the corresponding instruction, and were asked a binary question: “Evaluate whether the edited image successfully follows the given instruction.”
Appendix D VLMs failing to judge visual quality
You are an expert at evaluating image visual quality and naturalness.
I will show you an image.
1. Is the image visually pleasing?
2. Is the image natural?
3. Does the image look natural and coherent?
100 means the image is visually pleasing and natural.
0 means the image is not visually pleasing and natural.
50 means the image is neutral.
(a) Visual quality (Qwen2.5-VL: 60, HPSv3: 1.61).
(b) Visual quality (Qwen2.5-VL: 60, HPSv3: 3.35).
Figure 12: Examples of visual quality evaluation. Scores are reported from both VLM Qwen2.5-VL and HPSv3 for comparison. The VLM score ranges from 0 to 100, so the VLM gives a positive score to both images with significant artifacts.
Appendix E More Quality Examples
@@
Figure 13: T1: Change the color of pumpkin to purple; T2: Change the background to forest; T3: Remove fabric orange bow. Row-wise quality examples for the first four models: Nano Banana, GPT-Image-1, Gemini 2.0 Flash, and Qwen-Image-Edit. Each row shows generations for Input and three editing turns.

### v2

@@ Appendix G Discussion on Visual Quality
Low-level image statistics
@@
Appendix H VLMs failing to judge visual quality
You are an expert at evaluating image visual quality and naturalness.
I will show you an image.
1. Is the image visually pleasing?
2. Is the image natural?
3. Does the image look natural and coherent?
100 means the image is visually pleasing and natural.
0 means the image is not visually pleasing and natural.
50 means the image is neutral.
Appendix I Algorithmic Details
@@ I.1 Decomposition
  • Object listing: We use GPT-4o with the prompt in Section I.2.1. The model returns a JSON with one entry per object and a terminal aggregated string key ‘‘All Objects’’.
@@ I.2 Instruction Generation
8:  if not feasible($\tau_{t},\mathcal{O}$) then $\tau_{t}\leftarrow\texttt{subject\_add}$
@@
10:  $F_{t}\leftarrow\text{format\_instruction}(\tau_{t},\mathcal{O})$ $\triangleright$ Query VLM by prompts in Section I.2.2
@@
I.2.1 Object Listing Prompt
@@ I.3 Evaluation
We report HPSv3 Ma et al. (2025) plausibility and aesthetics, plus luminance metrics. Quality is not folded into the Overall score.
@@
I.4 Overall Score and Aggregation Details
@@ I.5 Model generations
Prompts are provided in full (Section I.2); thresholds are specified above. Grounding uses SwinT-OGC weights; consistency uses DINOv3 ViT-B/16; the quality head follows our RAHF implementation, and HPSv3 is included when available. All other parameters are left at defaults.
Appendix J Additional Evaluation Results
@@
For the multi-turn editing mode, the overall instruction-following success rate is reported in Table 8, while success rates for individual instruction types appear in Tables 9 and 10. Consistency results are summarized in Table 14. We also observed that some input images are non-square after resizing, which can leave black padding on the top/bottom or left/right edges. Certain editing models, such as GPT-Image-1 and Qwen-Image-Edit, attempt to fill these areas, whereas others preserve them. To account for this, we separately report consistency for square (Table 15) and non-square inputs (Table 16). The conclusions remain consistent with the overall evaluation. Quality results for multi-turn editing are presented in Table 18.
For the complex editing mode, the overall instruction-following success rate is shown in Table 11, and per-instruction-type results are in Tables 12 and 13. Consistency and quality results are reported in Tables 17 and 19, respectively.
@@
Seedream 4.0 75.93 55.58 41.59 75.93 75.58 76.11
Nano Banana 70.70 50.66 35.35 70.70 72.59 68.24
GPT-Image-1 73.12 54.89 38.35 73.12 74.44 73.12
FLUX.1-Kontext-max 69.49 46.89 31.83 69.49 69.11 70.43
Gemini 2.0 Flash 68.07 45.96 28.42 68.07 67.72 68.42
Qwen-Image-Edit 72.90 44.06 22.55 72.90 62.94 56.12
Step1X-Edit 61.89 34.97 17.83 61.89 59.09 53.32
FLUX.1-Kontext-dev 59.97 32.69 16.61 59.97 56.29 51.40
@@
Seedream 4.0 89.04 85.39 80.74 95.76 92.43 88.27 91.66 88.02 84.43 95.34 92.45 89.69
Nano Banana 88.68 85.19 83.31 96.88 93.79 92.41 89.73 87.64 88.48 94.21 93.77 93.36
GPT-Image-1 73.62 71.27 67.27 90.00 85.72 82.16 80.33 78.80 79.88 85.69 81.86 80.85
FLUX.1-Kontext-max 91.34 89.90 86.04 95.99 94.18 90.10 93.81 92.80 90.58 97.27 95.77 93.94
Gemini 2.0 Flash 84.16 79.91 71.13 91.50 91.83 87.07 88.98 87.07 85.13 95.19 93.35 93.34
Qwen-Image-Edit 75.36 72.68 68.85 88.50 88.98 82.71 81.66 78.47 77.04 91.16 87.85 85.78
Step1X-Edit 87.40 84.03 78.27 97.43 92.42 88.15 93.14 91.57 89.40 98.10 97.12 96.01
FLUX.1-Kontext-dev 92.55 88.92 84.04 96.49 93.57 92.19 92.83 90.61 88.34 96.91 95.74 94.73
@@
Seedream 4.0 89.55 83.48 81.23 95.49 92.38 90.86 93.50 88.85 86.44 94.17 88.62 84.35
Nano Banana 90.34 85.26 84.50 97.74 95.94 94.88 92.69 90.29 89.17 96.16 94.81 93.95
GPT-Image-1 73.22 68.53 67.20 88.60 86.88 83.99 79.57 78.27 77.35 77.20 75.91 75.14
FLUX.1-Kontext-max 90.86 86.31 83.43 97.06 95.27 93.51 94.20 91.09 89.34 95.73 93.16 91.14
Gemini 2.0 Flash 85.69 76.79 72.12 96.14 93.23 90.10 90.91 86.23 83.99 95.02 93.12 91.51
Qwen-Image-Edit 77.32 71.43 68.47 91.65 89.53 88.10 83.68 79.23 76.23 83.78 80.37 77.43
Step1X-Edit 88.26 81.37 77.22 97.33 95.77 93.77 94.01 90.53 88.73 98.26 97.09 95.70
FLUX.1-Kontext-dev 92.67 87.80 85.44 98.15 96.92 95.54 94.57 91.70 90.10 96.30 94.98 94.05
@@
Appendix K Human Agreement
@@
We conducted human study on edits made by four exemplary models—Step1X-Edit, AnyEdit, Gemini-Flash 2.0, and Flux.1-Kontext-dev—on EdiVal-Bench, generated by EdiVal-Agent as described in Section I.2. For each edit, we collected two human ratings, yielding a total of $572\times 4\times 2=4{,}576$ annotations. Depending on the prompt (which affected the editing instruction), each annotation took about 1–2 minutes. Raters were recruited online, each holding at least a bachelor’s degree. They were shown the original image, the edited image, and the corresponding instruction, and were asked a binary question: “Evaluate whether the edited image successfully follows the given instruction.”
Appendix L Counting
Among all subtasks, count_change is the most challenging. Even the best-performing model (GPT-Image-1) achieves a success rate below 25% at turn 1, while most models remain under 5%. We also provide illustrative examples in Figure 15.
(a) Base Image
(b) GPT-Image-1 (✓)
(c) Nano Banana (✓)
(d) Gemini 2.0 Flash (×)
(e) Qwen-Image-Edit (×)
Appendix M More Quality Examples
@@
Seedream 4.0 Refer to caption Refer to caption Refer to caption Refer to caption
@@
FLUX.1-Kontext-max Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 16: T1: Change the color of pumpkin to purple; T2: Change the background to forest; T3: Remove fabric orange bow. Row-wise quality examples for the first six models: Seedream 4.0, Nano Banana, GPT-Image-1, Gemini 2.0 Flash, FLUX.1-Kontext-max, and Qwen-Image-Edit. Each row shows generations for Input and three editing turns.